# Intrinsic training dynamics of deep neural networks

### v1

@@ 2.2 Conservation laws
Example˜2.5 illustrates a phenomenon that we will systematically exploit in our analysis: with the typical reparameterizations $\phi$ mentioned above, there exists a vector-valued function $\mathbf{h}:\theta\mapsto\mathbf{h}(\theta)\in\mathbb{R}^{N}$ that is conserved along the trajectory and allows to exhibit a function $K_{\theta_{0}}$ such that $M(\theta(t))=K_{\theta_{0}}(z(t))$ along the trajectory. As these will play a key role in our analysis we now introduce the essential concepts related to conservation laws.
We denote $\phi_{1},\cdots,\phi_{d}$ the $d$ coordinate functions of the reparameterization $\phi:\mathbb{R}^{D}\mapsto\phi(\theta)\in\mathbb{R}^{d}\in\mathcal{C}^{\infty}(\mathbb{R}^{D},\mathbb{R}^{d})$. Since $\phi$ yields a factorization of the loss, functions $h$ such that $\nabla h(\theta)\perp\nabla\phi_{i}(\theta)$ for each $i$ and each $\theta$ remain constant along the trajectory. This has been thoroughly analyzed, see e.g. Marcotte et al. (2023, 2024), using the following definition.
@@
If $\phi$ satisfies ˜2.1. and $h\in\mathcal{C}^{1}(\Omega,\mathbb{R})$ satisfies $\partial\phi(\theta)\nabla h(\theta)=0$ for every $\theta\in{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\Omega}$, then $h$ remains constant during the trajectory $\theta(t)$ of (1) for any initialization $\theta_{0}\in\Omega$.
The conservation laws associated with a given parameterization $\phi$ have been almost exhaustively studied for parameterizations corresponding to linear networks, ReLU networks, and attention layers. In particular, prior work has shown that all conservation laws for $\phi$ in the cases of ReLU (cf Example˜2.3) and linear (cf Example˜2.2) networks (see Marcotte et al. (2023)) as well as for an attention layer (see Marcotte et al. ) are captured by the following proposition (Marcotte et al. (2023)). This has been proven theoretically for two-layer networks and empirically validated for deeper architectures using symbolic computation (see Marcotte et al. (2023)). It is worth noticing that all conservation laws in such cases are polynomials.
@@
are conservation laws for $\phi_{\mathtt{Lin}}$ (resp. $\phi_{\mathtt{ReLU}}$). Similary, considering $\theta\coloneqq(Q,K,V,O)$ and $\phi_{\mathtt{Att}}$ from Example˜2.4, $\mathbf{h}:\theta\mapsto(QQ^{\top}-KK^{\top},VV^{\top}-OO^{\top})$ is a set of conservation laws for $\phi_{\mathtt{Att}}$.
2.2.1 Intrinsic dynamics via conservation laws
Given conservation laws $\mathbf{h}(\theta)$ for $\phi$, the trajectory $\theta(t)$ for (1) remains at all times on the set
@@ 2.2.1 Intrinsic dynamics via conservation laws
determined by the initialization $\theta_{0}$. This holds true for any function $f$ such that $\ell=f\circ\phi$ (hence, in machine learning: for any task/loss and any dataset, provided that the neural network model is (locally) factorized via $\phi$).
To establish the existence of a function $K_{\theta}(\cdot)$ such that $M(\theta(t))=K_{\theta_{0}}(z(t))$ on the whole trajectory, a natural relaxation is thus to establish a related equality on the whole set $\mathcal{M}_{\theta_{0}}$ rather than only on a specific trajectory. This leads to the following definition and its immediate consequence.
@@
We say that $\theta_{0}$ verifies the intrinsic metric property on an open set $U\ni\theta_{0}$ with respect to $\phi$, if there exists conservation laws $\mathbf{h}(\theta)\in\mathbb{R}^{N}$ for $\phi$ and a function $K_{\theta_{0}}\in\mathcal{C}^{1}(\mathbb{R}^{d},\mathbb{R}^{d\times d})$ such that $M(\theta)=K_{\theta_{0}}(\phi(\theta))$ for each $\theta\in\mathcal{M}_{\theta_{0}}\cap U$.
@@
Lemma˜2.11 remains valid with a slightly weakened version of Definition˜2.10, where $K_{\theta_{0}}$ is not required to be smooth. Yet, since the existence of a smooth solution to the resulting intrinsic ODE (5) is simplified when $K_{\theta_{0}}$ is $\mathcal{C}^{1}$ we chose to include this in the definition.
The following theorem (proved in Appendix˜B) establishes a necessary condition for the intrinsic metric property to hold. We use it to show that the property does not always hold for linear networks.
@@
    There exists an open set $O\supset\phi(\mathcal{M}_{\theta_{0}})\cap U$ and a map $K_{\theta_{0}}\in\mathcal{C}^{1}(O,\mathbb{R}^{d\times d})$ such that for each $\theta\in\mathcal{M}_{\theta_{0}}\cap U$: $M(\theta)=K_{\theta_{0}}(\phi(\theta))$;
@@
    For every $\theta\in\mathcal{M}_{\theta_{0}}\cap U$ one has
    $\mathrm{ker}\partial\phi(\theta)\cap\mathrm{ker}\partial\mathbf{h}(\theta)\subseteq\mathrm{ker}\partial M(\theta).$ (7)
On one hand, a trivial case where (7) holds is when the intersection of kernels on the left hand side is zero:
@@
This stronger assumption can in fact be shown to imply the intrinsic metric property (see Theorem˜3.3 in the upcoming section), and we will show (cf Corollary˜3.9) that, with $\phi_{\mathtt{ReLU}}$ associated to general ReLU networks of any depth, there exists a set of conservation laws such that (8) indeed holds for any initialization. This implies the intrinsic metric property and therefore the intrinsic dynamic property irrespective of the initialization for ReLU networks with $\phi_{\mathtt{ReLU}}$.
On the other hand, for linear networks with more than one hidden neuron, we will show that it is not possible to reduce the problem to (8). Nevertheless, certain initializations (known as balanced conditions (Arora et al. (2019)) are known to satisfy the intrinsic metric property with respect to the reparametrisation $\phi_{\mathtt{Lin}}$ (cf (Arora et al., 2018, Theorem 1), (Bah et al., 2022, Lemma 2)). In this paper, we generalize this result to so-called relaxed balanced initializations (see Definition˜4.1). Moreover, we show that in certain configurations, relaxed balanced initializations are exactly the only ones that satisfy the intrinsic metric property (cf Theorem˜4.3 and Theorem˜4.4).
3 Intrinsic recoverability and application for ReLU neural networks
In this section we consider a stronger condition called intrinsic recoverability property which requires not only that $M(\theta(t))$ can be rewritten as a function of $z(t)$ and the initialization, but that at each point of the trajectory $\theta(t)$ itself can be fully expressed in terms of $z(t)$ and the initialization $\theta_{0}$. In other words, in this scenario, $\theta(t)$ can be completely recovered from the parameterization $\phi$ and the initialization alone, hence the name. As we will establish, this apparently strong property indeed holds when (8) is satisfied, which is always the case for ReLU networks.
3.1 Intrinsic recoverability implies intrinsic metric
@@
The intrinsic recoverability property is equivalent to (8) (see Appendix˜C for a proof):
@@
Consider $\phi\in\mathcal{C}^{2}(\mathbb{R}^{D},\mathbb{R}^{d})$ and $\theta_{0}\in\mathbb{R}^{D}$. The following are equivalent: (i) there exists conservation laws $\mathbf{h}\in\mathcal{C}^{1}(\Omega,\mathbb{R}^{N})$ for $\phi$ on a neighborhood $\Omega$ of $\theta_{0}$ such that (8) holds for each $\theta\in\mathcal{M}_{\theta_{0}}\cap\Omega$; (ii) there exists an open set $U\subseteq\Omega$ on which $\theta_{0}$ satisfies the intrinsic recoverability property Definition˜3.1 (and thus the intrinsic metric property Definition˜2.10) with respect to $\phi$.
3.2 The Frobenius property is almost equivalent to intrinsic recoverability
We are interested in establishing condition (8), as it implies the intrinsic recoverability property, and thus an intrinsic dynamics. It may not seem obvious a priori how to verify whether such a condition can hold, nor how to construct suitable conservation laws $\mathbf{h}$ in practice. Intuitively, one should select as many conservation laws as possible while ensuring they remain independent, in a specific sense defined by (Marcotte et al., 2023, Definition 2.18). As shown by Marcotte et al. (2023), knowing the maximal number of such conservation laws can be checked using Lie brackets of the associated vector fields. We recall the relevant definitions and explain how this criterion applies in our setting.
@@ 3.2 The Frobenius property is almost equivalent to intrinsic recoverability
Given two vector fields $\chi_{1},\chi_{2}\in\mathcal{C}^{\infty}(\Theta,\mathbb{R}^{d})$, the Lie brackets $\chi_{1},\chi_{2}]$ is the vector field defined by $\chi_{1},\chi_{2}](\theta)\coloneqq\partial\chi_{2}(\theta)\chi_{1}(\theta)-\partial\chi_{1}(\theta)\chi_{2}(\theta)$.
@@
Given some function space $\mathbb{W}\subseteq\mathcal{C}^{\infty}(\Theta,\mathbb{R}^{d})$, the generated Lie algebra of $\mathbb{W}$ is the smallest subspace of $\mathcal{C}^{\infty}(\Theta,\mathbb{R}^{d})$ that contains $\mathbb{W}$ and that is stable by Lie brackets, and is denoted $\mathrm{Lie}(\mathbb{W})$.
@@
and for any infinitly smooth $\phi$ we denote $\mathbb{W}_{\phi}\coloneqq\mathrm{span}\{\nabla\phi_{i}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(\cdot),\ 1\leq i\leq d}\}\subseteq\mathcal{C}^{\infty}(\Theta,\mathbb{R}^{d})$.
@@
An infinitely smooth function $\phi$ satisfies the Frobenius property on $\Omega$ if for all $\theta\in\Omega$, $\mathrm{Lie}(\mathbb{W}_{\phi})(\theta)=\mathbb{W}_{\phi}(\theta)$.
The following proposition (see Appendix˜D for a proof) shows how this property implies the intrinsic dynamic property of $\theta_{0}$. In particular, as Frobenius property does not hold for $\phi_{\mathtt{Lin}}$ (Marcotte et al., 2023, Proposition I.1), it is not possible to have the intrinsic recoverability for linear networks with classical $\phi_{\mathtt{Lin}}$.
@@
We have the following implications $\implies(ii)\implies(iii): (i)(i)$ $\phi$ satisfies the Frobenius property on $\Omega$ and the trace of $\mathbb{W}_{\phi}$ has its dimension that is constant on $\Omega$; $(ii)$ For any $\theta_{0}\in\Omega$, there exists conservation laws $\mathbf{h}$ for $\phi$ on a neighborhood $U\subset\Omega$ of $\theta_{0}$ such that for each $\theta\in\mathcal{M}_{\theta_{0}}$ (8) holds; $(iii)$ $\phi$ satisfies the Frobenius property on $\Omega$.
The result presented in Li et al. (2022) can be recovered as a special case of Proposition˜3.7: the authors require the reparametrization $\phi$ to be commuting, meaning that for all pairs $\phi_{i},\phi_{j}$, the Lie bracket $\nabla\phi_{i},\nabla\phi_{j}]$ is equal to zero. In this setting, $\phi$ naturally satisfies the Frobenius property, and the result of Li et al. (2022) establishes an even stronger property: the dynamics on $\phi(\theta)$ form a mirror flow. In particular, it is worth noting that diagonal networks satisfy that their parametrization (product of the diagonals) is commuting (as all coordinates functions are separable), which thus (Li et al. (2022)) implies a mirror flow dynamic, and thus an implicit bias (see e.g. Azulay et al. (2021)). In contrast, we consider a weaker condition; we seek only to determine whether the dynamics on $z=\phi(\theta)$ can be expressed intrinsically as a Riemannian gradient flow.
3.3 Application to general deep ReLU networks
We now show that the intrinsic recoverability property is satisfied for any initialisation for the parametrisation $\phi$ associated to a large class of (deep) ReLU neural networks. While this result is already known in the two-layer case (Marcotte et al., 2023, Examples 3.5 and 3.8), here we establish it for the general model of ReLU networks of Gonon et al. (2023), associated to a directed acyclic graph (DAG) of any depth, including skip connections and arbitrary mixes of ReLU/linear/max-pooling activations. We first establish that $\phi_{\mathtt{ReLU}}$ satisfies the Frobenius property (see Appendix˜E for a proof):

### v2

@@ 2.2 Conservation laws
Example˜2.5 illustrates a phenomenon that we will systematically exploit in our analysis: with the typical reparameterizations $\phi$ mentioned above, there exists a vector-valued function $\mathbf{h}:\theta\mapsto\mathbf{h}(\theta)\in\mathbb{R}^{N}$ that is conserved along the trajectory and allows to exhibit a function $K_{\theta_{0}}$ such that $M(\theta(t))=K_{\theta_{0}}(z(t))$ along the trajectory. As these will play a key role in our analysis we now remind the essential concepts related to conservation laws.
We denote $\phi_{1},\cdots,\phi_{d}$ the $d$ coordinate functions of the reparameterization $\phi:\mathbb{R}^{D}\mapsto\phi(\theta)\in\mathbb{R}^{d}\in\mathcal{C}^{\infty}(\mathbb{R}^{D},\mathbb{R}^{d})$. Since $\phi$ yields a factorization of the loss, functions $h$ such that $\nabla h(\theta)\perp\nabla\phi_{i}(\theta)$ for each $i$ and each $\theta$ remain constant along the trajectory. This has been thoroughly analyzed, see e.g. Marcotte et al. (2023; 2024), using the following definition.
@@
[…]
The conservation laws associated with a given parameterization $\phi$ have been almost exhaustively studied for parameterizations corresponding to linear networks, ReLU networks, and attention layers. In particular, prior work has shown that all conservation laws for $\phi$ in the cases of ReLU (cf Example˜2.3) and linear (cf Example˜2.2) networks (see Marcotte et al. (2023)) as well as for an attention layer (see Marcotte et al. (2025)) are captured by the following proposition (Marcotte et al., 2023). In Le and Jegelka (2022), the authors also give canonical conservation laws for general DAG ReLU architectures. We recall these conservation laws in Lemma˜F.9. This has been proven theoretically for two-layer networks and empirically validated for deeper architectures using symbolic computation (see Marcotte et al. (2023)). One of the contributions of this paper (see Corollary˜3.4 and Corollary˜F.15) is to establish the completeness of conservation laws for $\phi$ associated with any DAG ReLU network. It is worth noticing that all conservation laws in such cases are polynomials.
@@
are conservation laws for $\phi_{\mathtt{Lin}}$ (resp. $\phi_{\mathtt{ReLU}}$). Similarly, considering $\theta\coloneqq(Q,K,V,O)$ and $\phi_{\mathtt{Att}}$ from Example˜2.4, $\mathbf{h}:\theta\mapsto(QQ^{\top}-KK^{\top},VV^{\top}-OO^{\top})$ is a set of conservation laws for $\phi_{\mathtt{Att}}$.
2.3 Intrinsic dynamics via conservation laws
Given conservation laws $\mathbf{h}(\theta)$ for $\phi$, any trajectory $\theta(t)$ satisfying equation 1 remains at all times on the set
@@ 2.3 Intrinsic dynamics via conservation laws
determined by $\theta_{0}$. This holds true for any function $f$ such that $\ell=f\circ\phi$ (hence, in machine learning: for any task/loss and any dataset, provided that the network model is (locally) factorized via $\phi$).
To establish the existence of a function $K_{\theta_{0}}(\cdot)$ such that $M(\theta(t))=K_{\theta_{0}}(z(t))$ on the whole trajectory, a natural relaxation is thus to establish a related equality on the whole set $\mathcal{M}_{\theta_{0}}$ rather than only on a specific trajectory. This leads to the following definition and its immediate consequence.
@@
We say that $\theta_{0}$ verifies the intrinsic metric property with respect to $\phi$ on an open set $U\ni\theta_{0}$ if there exists conservation laws $\mathbf{h}(\theta)\in\mathbb{R}^{N}$ for $\phi$ and a function $K_{\theta_{0}}\in\mathcal{C}^{1}(\mathbb{R}^{d},\mathbb{R}^{d\times d})$ such that $M(\theta)=K_{\theta_{0}}(\phi(\theta))$ for each $\theta\in\mathcal{M}_{\theta_{0}}\cap U$.
@@
Lemma˜2.11 remains valid with a slightly weakened version of Definition˜2.10, where $K_{\theta_{0}}$ is not required to be smooth. Yet, since the existence of a smooth solution to the resulting intrinsic ODE equation 5 is simplified when $K_{\theta_{0}}$ is $\mathcal{C}^{1}$, we chose to include this in the definition.
The following theorem (proved in Appendix˜C) establishes a necessary condition for the intrinsic metric property to hold. We use it to show that the property does not always hold for linear networks.
@@
    There exists an open set $O\supset\phi(\mathcal{M}_{\theta_{0}}\cap U)$ and a map $K_{\theta_{0}}\in\mathcal{C}^{1}(O,\mathbb{R}^{d\times d})$ such that:
    $M(\theta)$ $=K_{\theta_{0}}(\phi(\theta)),\qquad\forall\theta\in\mathcal{M}_{\theta_{0}}\cap U;$
@@
    $\mathrm{ker}\partial\phi(\theta)\cap\mathrm{ker}\partial\mathbf{h}(\theta)\subseteq\mathrm{ker}\partial M(\theta),\quad\forall\theta\in\mathcal{M}_{\theta_{0}}\cap U$. (7)
A trivial case where equation (ii) holds is when the intersection of kernels on the left hand side is zero:
@@
This stronger assumption can in fact be shown to imply the intrinsic metric property (see Theorem˜2.17 in the upcoming section), and we will show (cf Corollary˜3.3) that, with $\phi_{\mathtt{ReLU}}$ associated to general DAG (directed acyclic graph) ReLU networks of any depth, there exists a set of conservation laws such that equation 8 indeed holds for any initialization. This implies the intrinsic metric property and therefore the intrinsic dynamic property irrespective of the initialization for ReLU networks with $\phi_{\mathtt{ReLU}}$.
For linear networks with more than one hidden neuron, equation 8 does not hold (as a consequence of Steps 2,3 in the proof of Theorem˜3.9), and we will show that equation (ii) does not always hold. Nevertheless, certain initializations, known as balanced conditions (Arora et al., 2019), are known to satisfy the intrinsic dynamic property with respect to the reparametrisation $\phi_{\mathtt{Lin}}$ (cf (Arora et al., 2018, Theorem 1), (Bah et al., 2022, Lemma 2)). In this paper, we generalize this result to so-called relaxed balanced initializations (see Definition˜3.6). Moreover, we show that in certain configurations, relaxed balanced initializations are exactly the only ones that satisfy the intrinsic metric property (cf Theorem˜3.8 and Theorem˜3.9).
2.4 Intrinsic recoverability
In this section we consider a stronger condition called intrinsic recoverability property which requires not only that $M(\theta(t))$ can be rewritten as a function of $z(t)$ and the initialization, but also that at each point of the trajectory $\theta(t)$ itself can be fully expressed in terms of $z(t)$ and the initialization $\theta_{0}$. In other words, in this scenario, $\theta(t)$ can be completely recovered from the parameterization $\phi$ and the initialization alone, hence the name. As we will establish, this apparently strong property indeed holds when equation 8 is satisfied, which is always the case for ReLU networks.
2.4.1 Intrinsic recoverability implies intrinsic metric
@@
The intrinsic recoverability property is equivalent to equation 8 (see Appendix˜D for a proof):
@@
Given $\phi\in\mathcal{C}^{2}(\mathbb{R}^{D},\mathbb{R}^{d})$ and $\theta_{0}\in\mathbb{R}^{D}$, the following are equivalent: (i) there are conservation laws $\mathbf{h}\in\mathcal{C}^{1}(\Omega,\mathbb{R}^{N})$ for $\phi$ on a neighborhood $\Omega$ of $\theta_{0}$ such that equation 8 holds for each $\theta\in\mathcal{M}_{\theta_{0}}\cap\Omega$; (ii) there is an open set $U\subseteq\Omega$ on which $\theta_{0}$ satisfies the intrinsic recoverability property Definition˜2.15 (and thus the intrinsic metric property Definition˜2.10) with respect to $\phi$.
2.4.2 The Frobenius property is almost equivalent to intrinsic recoverability
We are interested in equation 8, as it implies the intrinsic recoverability property, and thus an intrinsic dynamic. It may not seem obvious a priori how to verify whether such a condition can hold, nor how to construct suitable conservation laws $\mathbf{h}$ in practice. Intuitively, one should select as many conservation laws as possible while ensuring they remain independent, in a specific sense (Marcotte et al., 2023, Definition 2.18). As shown by Marcotte et al. (2023), knowing the maximal number of such conservation laws can be checked using Lie brackets of the associated vector fields. We recall the relevant definitions and explain how this criterion applies in our setting.
@@ 2.4.2 The Frobenius property is almost equivalent to intrinsic recoverability
Given two vector fields $\chi_{1},\chi_{2}\in\mathcal{C}^{\infty}(\Theta,\mathbb{R}^{d})$, the Lie bracket $\chi_{1},\chi_{2}]$ is the vector field defined by $\chi_{1},\chi_{2}](\theta)\coloneqq\partial\chi_{2}(\theta)\chi_{1}(\theta)-\partial\chi_{1}(\theta)\chi_{2}(\theta)$.
@@
Given some function space $\mathbb{W}\subseteq\mathcal{C}^{\infty}(\Theta,\mathbb{R}^{d})$, the generated Lie algebra of $\mathbb{W}$ is the smallest subspace of $\mathcal{C}^{\infty}(\Theta,\mathbb{R}^{d})$ that contains $\mathbb{W}$ and that is stable under Lie brackets, and is denoted $\mathrm{Lie}(\mathbb{W})$.
@@
and for any infinitely smooth $\phi$ we denote $\mathbb{W}_{\phi}\coloneqq\mathrm{span}\{\nabla\phi_{i}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(\cdot),\ 1\leq i\leq d}\}\subseteq\mathcal{C}^{\infty}(\Theta,\mathbb{R}^{D})$.
@@
A $\mathcal{C}^{\infty}$ function $\phi$ satisfies the Frobenius property on $\Omega$ if for all $\theta\in\Omega$, $\mathrm{Lie}(\mathbb{W}_{\phi})(\theta)=\mathbb{W}_{\phi}(\theta)$. This property is slightly weaker than involutivity (Isidori, 1995).
The following proposition (proved in Appendix˜E) relates this property to the intrinsic dynamic property of $\theta_{0}$. In particular, as the Frobenius property does not hold for $\phi_{\mathtt{Lin}}$ (Marcotte et al., 2023, Proposition I.1), it is not possible to have the intrinsic recoverability for linear networks with classical $\phi_{\mathtt{Lin}}$.
@@
We have the following implications $\implies(ii)\implies(iii): (i)(i)$ $\phi$ satisfies the Frobenius property on $\Omega$ and the trace of $\mathbb{W}_{\phi}$ has its dimension that is constant on $\Omega$; $(ii)$ For any $\theta_{0}\in\Omega$, there exists conservation laws $\mathbf{h}$ for $\phi$ on a neighborhood $U\subset\Omega$ of $\theta_{0}$ such that for each $\theta\in\mathcal{M}_{\theta_{0}}$ equation 8 holds; $(iii)$ $\phi$ satisfies the Frobenius property on $\Omega$.
Link with the commuting property. The result presented in Li et al. (2022) can be recovered as a special case of Proposition˜2.21: the authors require the reparametrization $\phi$ to be commuting, meaning that for all pairs $\phi_{i},\phi_{j}$, the Lie bracket $\nabla\phi_{i},\nabla\phi_{j}]$ is zero. In this setting, $\phi$ naturally satisfies the Frobenius property, and the result of Li et al. (2022) establishes an even stronger property: the dynamics on $\phi(\theta)$ form a mirror flow. In particular, it is worth noting that diagonal networks satisfy that their parametrization (product of the diagonals) is commuting (as all coordinates functions are separable), which thus (Li et al., 2022) implies a mirror flow dynamic, and thus an implicit bias (see e.g. Azulay et al. (2021)). In contrast, we consider a weaker condition; we seek only to determine whether the dynamics on $z=\phi(\theta)$ can be expressed intrinsically as a Riemannian gradient flow.
3 Application for general ReLU networks and Linear Networks
We now show that the intrinsic recoverability property is satisfied for any initialization for the parametrization $\phi$ associated to a large class of (deep) ReLU neural networks. While this result is already known in the two-layer case (Marcotte et al., 2023, Examples 3.5 and 3.8), here we establish it for the general model of ReLU networks of Gonon et al. (2024), associated to a directed acyclic graph (DAG) of any depth, including skip connexions and arbitrary mixes of ReLU/linear/max-pooling activations. We first establish that $\phi_{\mathtt{ReLU}}$ satisfies the Frobenius property (see Section˜F.1 for a proof):